Friday 09/07/12

9:05 A – 8:00 P

  • O/N STORM lost focus
  • attempt to get bead images aborted.  Issues with focus lock delayed start and Colenso needs to get on the machine.
  • O/N confocal still had lots of issues with z-focus (apparently not all embryos positions were set to same plane?)   Some still look usable.  Downloaded data.
  • O/N meta-periodiate treatment:  en samples detached from slide.  much less bright (less autofluorescence and signal?).  Not eched, resin blocks clearly still visible.
  • centromere-STORM coverslip remained attached.  samples might be thinner – but still in resin.  seems to uniformly digest away resin and sample equally (and very slowly).
  • set up Pc-GFP data to run on Tuck (may take rest of weekend?)
  • Attempt to download ML slide data
  • Update from Jacques: also modeling competitive repression for promoter paper. This appears to conflict with bistability paper modeling.  Should be discussed.  sent email to Mike.
  • Mufit python: parameter ‘cycles’ controls the number of iterations and affects the number of potential overlapping molecules that can be distinguished.  For python implementation of single emitter fitting (inlcuding big blunt objects) set this parameter to 1.  Should substantially speed up code.  Try this on Odyssey.
  • Working on a bistability manuscript.
This entry was posted in Summaries and tagged , . Bookmark the permalink.